OpenStudy (anonymous):

Which is the simplified form of 12 times a to the sixth power all over 8 times a cubed.

OpenStudy (anonymous):

\[\huge \frac{12a^6}{8a^3} =\frac{3a^{6-3}}{2} =\frac{3a^{3}}{2} \]

OpenStudy (mathstudent55):

Is this the question? \( \dfrac{12a^6}{8a^3}\)